Which number is one-fourth between 2.2 and 8.5?
I chose 3.775, I am correct?

Answer by Alan3354(69443)

Which number is one-fourth between 2.2 and 8.5?
I chose 3.775, I am correct?
============
Diff = 8.5 - 2.2 = 6.3
6.3/4 = 1.575
2.2 + 1.575 = 3.775
------------
8.5 - 1.575 is also "1/4 between"
Not a well-worded question.